Find a pair of numbers with a sum of 26 and a product that is a maximum. Find the maximum product.
Answer by oberobic(2304)

x + y = 26
so
x = 26-y
.
The product is
z = x * y = (26-y) * y
.
z = 26y -y^2
.
To find the maximum, take the first derivative and solve it.
.
dz/dy = 26 - 2y
.
26-2y = 0
26 = 2y
y = 13
.
Check:
y      x=26-y   x*y   
1313169
1214168
1115165
1016160
917153
818144
719133
620120
521105
42288
32369
22448
12525

.
Answer: The pair of numbers are 13 and 13 that total 26 and maximizes the product.
.
Done.